condense the expression log 3+ 1/2 log x-log 5
Found 2 solutions by chiefman, Ploy:
Answer by chiefman(11)

log3+1/2logx-log5 this becomes
log3+logx^1/2-log5 since( alogb=logb^a)
log{(3*x^1/2)/5}since[loga+logb=logab],and [loga-logb=log(a/b)]
hence the answer is
log{(3*x^1/2)/5}

Answer by Ploy(1)

log3+logx^1/2-log5 = 0
log(3/5)(x^1/2) = 0
log(3*x^1/2)/5 = 0
(3*x^1/2)/5 = 10^0
(3*x^1/2)/5 = 1
3*x^1/2 = 5
x^1/2 = 5/3
x = 25/9
x = 2.78